# Probabilistic Robustness for Free? Revisiting Training via a Benchmark

### v1

@@ 1 Introduction
Deep learning (DL) has demonstrated remarkable potential to drive transformative advancements across a wide range of industries, such as autonomous driving and medical diagnostics. In such safety-critical domains, robustness is a fundamental prerequisite for DL models’ pervasive deployment. Numerous studies have investigated DL robustness, leading to a range of benchmarks that systematically track the progress in the topic Ling et al. (2019); Guo et al. (2023); Croce et al. (2021); Tang et al. (2021); Dong et al. (2020); Liu et al. (2025). These efforts have predominantly focused on adversarial robustness (AR), which emphasizes the extreme worst-case scenarios by evaluating local robustness based on the existence of deterministic adversarial examples (AEs): subtle perturbations to inputs that alter model predictions Szegedy et al. (2014); Goodfellow (2015); Papernot et al. (2016b).
@@
A more recent and practical perspective, probabilistic robustness (PR) Webb et al. (2019); Weng et al. (2019); Couellan (2021); Baluta et al. (2021); Zhao et al. (2021); Tit et al. (2021); Robey et al. (2022); Pautov et al. (2022); Zhang et al. (2023b); Huang et al. (2023); Dong et al. (2023); TIT et al. (2023); Zhang et al. (2024a; b; 2025); Zhang & Sun (2025); Zhang et al. (2023a), employs statistical approaches to answer the question: “What is the probability that predictions remain correct under stochastic perturbations”. This probabilistic view is arguably more relevant to real-world applications than AR, as it provides an overall assessment of a model’s local robustness, accounting for scenarios where AEs may exist Webb et al. (2019); Huang et al. (2023); Zhao (2025) and acknowledging residual risks Zhang et al. (2023b; 2024a; 2025) that are more realistic to manage in practice.
@@
Despite its potential, most existing work on PR is restricted to evaluation, with only a few studies Wang et al. (2021); Robey et al. (2022); Zhang et al. (2024a; 2025) exploring training methods specifically designed to improve PR Zhao (2025). Among them, we identify three limitations. First, they use different sets of evaluation metrics, with none adopting a comprehensive set to assess the training methods holistically, making it difficult to understand which methods are truly effective/efficient and under what conditions. Second, an interesting observation is that adversarial training (AT) that primarily designed to improve AR often enhances PR as a “free by-product”. However, existing studies assess only a limited selection of AT methods, limiting the understanding of AT’s potential as a more efficient PR improvement tool. Third, there is no unified theoretical framework to compare the generalisability of PR-targeted training, making it unclear of their broader applicability in various context. These gaps and the rapid growth of PR research highlight the need for a systematic benchmark for PR training methods, which is currently missing from the literature.
@@
To bridge the gaps, we introduce $\mathtt{PRBench}$, the first benchmark dedicated to evaluating training methods for improving PR. Key features of $\mathtt{PRBench}$ include: i) A comprehensive set of metrics, covering clean accuracy, PR and AR performance, training efficiency, and generalisation error (GE); ii) A large set of training methods including most common AT methods and all PR-targeted training methods; iii) Theoretical bounds of GE are derived under a unified framework of Uniform Stability Analysis Xiao et al. (2022b); Cheng et al. to conclude comparative insights. Specifically, $\mathtt{PRBench}$ includes 222 trained models based on 7 widely adopted datasets and 10 model architectures. It uses 4 common adversarial attacks to measure AR performance, 2 types of PR metrics (with various hyperparameters, e.g., distributions of perturbations) for PR, GE metrics for both AR and PR, the clean accuracy and the training time. While the total number of training methods included is 13, $\mathtt{PRBench}$ code-base is designed to be extendable for future inclusion and comparison of new methods.
@@ 2.1 Generalisation Errors and Adversarial Robustness
AT Goodfellow (2015); Tramèr et al. (2018); Gowal et al. (2019); Madry et al. (2018); Balunović & Vechev (2020) is the most common and effective empirical approach for enhancing AR. It is typically formulated as a min-max optimisation problem, where the inner max aims at finding the worst-case AE in Eq. 2, e.g., by FGSM Goodfellow (2015) and PGD Madry et al. (2018). Then the model is trained to minimise the loss over these AEs derived from a training dataset $\mathcal{D}$:
@@ 2.2 Probabilistic Robustness
PR adopts a probabilistic view, evaluating the overall local robustness in the presence of AEs, cf. Fig. 1 (b). This probabilistic notion of robustness is acknowledged to be more practical than the worst-case AR, as many applications only require the risk of AEs to stay below an acceptable threshold, rather than eliminating them entirely Webb et al. (2019); Zhang et al. (2023b; 2024b); Zhao (2025).
@@
where $I_{\mathcal{S}}(\bm{x})$ is an indicator function that equals 1 when $\mathcal{S}$ is true and 0 otherwise; $Pr(\cdot)$ is the local distribution of inputs representing how perturbations $\bm{\delta}$ are generated, which is precisely the “input model” used by Webb et al. (2019); Weng et al. (2019); Zhang et al. (2024b).
Def. 1 suggests that PR is the probability that the model prediction remains unchanged from a random perturbation $\bm{x}^{\prime}$. A “frequentist” interpretation of this expected probability is—it is the limiting relative frequency of perturbations where the output label is preserved, in an infinite sequence of independently generated perturbations Zhang et al. (2024b). In other words, the “proportion” of non-AEs in the infinite set of perturbed inputs. To evaluate PR, one of the earliest works Webb et al. (2019) introduced a useful black-box statistical estimator, especially for cases where PR is very high. Later, more efficient white-box statistical estimators were proposed in TIT et al. (2023). Additionally, PR has been extended to applications such as explainable AI Huang et al. (2023) and text-to-image models Zhang et al. (2024b). For an overview of PR estimators, we refer readers to Zhao (2025).
In contrast to the relatively extensive studies on PR estimators, research on training methods for PR remains scarce. To the best of our knowledge, four studies Wang et al. (2021); Robey et al. (2022); Zhang et al. (2024a; 2025) explicitly motivated to develop training methods for improving PR. The first three can be categorized as Risk-based Training (RT) methods, which shift away from the worst-case perspective towards a training paradigm based on statistical risks induced by distributional perturbations, incorporating perturbation risk functions into the objective. We formally introduce its general formulation as:
@@
The specific choice of $\mathcal{R}_{\gamma}$ in RT methods varies: Wang et al. (2021) uses the identity function, while Robey et al. (2022); Zhang et al. (2024a) adopt functions calculating the Conditional Value-at-Risk (CVaR) and Entropic Value-at-Risk (EVaR), respectively. Essentially, rather than training on optimized AEs like AT methods for AR, RT methods train on statistical risks quantified by sampling stochastic perturbations from the perturbation distribution.
A very recent idea of adapting AT for PR was proposed in Zhang et al. (2025), which does not follow RT paradigm of Def. 2. Instead, their approach aligns with the traditional AT formulation in Eq. 3. Through multi-start PGD attacks and decision boundary exploration, AT-PR identifies an optimal AE whose neighborhood constitutes the largest all-AE region. Intuitively, training over such an optimal AE would reduce the overall “proportion” of AEs in the local norm ball and thus improve PR. We classify it as “hybrid” given it is targeting PR but adapting AT by solving a new min-max problem. Both RT methods and the “hybrid” method AT-PR constitute the current class of PR-targeted method. Further details for all four training methods are provided in Appendix C.
@@ 2.3 Robust Overfitting
While AT is regarded as one of the most promising methods for enhancing AR, empirical studies Rice et al. (2020); Gowal et al. (2020) have shown that it suffers significantly from overfitting. To understand this phenomenon, several theoretical studies have been conducted under different frameworks, including VC-dimension Montasser et al. (2019); Attias et al. (2022), Rademacher complexity Khim & Loh (2018); Yin et al. (2019); Awasthi et al. (2020); Xiao et al. (2022a), and Uniform Algorithmic Stability Farnia & Ozdaglar (2021); Xing et al. (2021); Xiao et al. (2022c). Among them, one of the most notable contributions is by Xiao et al. (2022b), which introduces the concept of $\eta$-approximate $\beta$-smoothness, as defined in Eq. 6. This work Xiao et al. (2022b) analyses the surrogate loss of the max function and show that generalisation is affected by an additional term proportional to $\eta$.
Definition 3 (Approximate Smoothness Xiao et al. (2022b))
@@
This assumption has been widely adopted and validated in prior work Farnia & Ozdaglar (2021); Xing et al. (2021); Xiao et al. (2022c; b), where it serves as a foundational premise for analyzing algorithmic stability and generalisation under AT. We show that the GE is bounded in Thm. 1.
@@
Thm. 1 follows from the results in Xiao et al. (2022b), where the authors conclude that the additional term $n\varphi\phi$ contributes to robust overfitting. The complete proof corresponds to the special case $\lambda=0$ of Thm. 3 in Appendix E.2. And the proof for uniform stability is shown in Appendix E.5.
@@ 3 Description of PRBench
All existing robustness benchmarks focus exclusively on AR Guo et al. (2023); Croce et al. (2021); Tang et al. (2021); Dong et al. (2020); Liu et al. (2025), with more related benchmarks summarized in Appendix A.5. In contrast, our $\mathtt{PRBench}$ is the first benchmark dedicated to PR. This section describes the models, datasets, training methods, and evaluation metrics in $\mathtt{PRBench}$.
@@ 3.2 Training Methods
For AT, we consider the following representative methods. The PGD attack is widely regarded as the standard approach, which minimizes the adversarial cross-entropy loss $\mathcal{L}_{\textit{CE}}$. Building on this, logit pairing methods Kannan et al. (2018), including adversarial logit pairing (ALP) and clean logit pairing (CLP), augment the objective with a regularization term coupling natural and AEs. TRADES further introduces a KL-based regularization, defining its objective as a combination of the natural loss and the KL divergence $\mathcal{L}_{\textit{KL}}$ between predictions on clean and adversarial inputs. MART follows the same KL-based regularization but augments it with the adversarial cross-entropy loss $\mathcal{L}_{\textit{CE}}$, while additionally emphasizing misclassified examples through larger penalty weights. Overall, existing AT methods can be characterized by different uses of $\mathcal{L}_{\textit{CE}}$ and $\mathcal{L}_{\textit{KL}}$, either individually or in combination, for both AE generation and training. We identify a missing configuration and introduce KL-PGD, a combination that generates AEs using $\mathcal{L}_{\textit{KL}}$ (as in TRADES) and trains the model with $\mathcal{L}_{\textit{CE}}$ (as in PGD). KL-PGD is designed as a diagnostic variant to disentangle the effects of loss functions from perturbation generation strategies, thereby providing a controlled comparison point for systematically evaluating the relative contributions of different optimization objectives.
@@
By instantiating the general formulation of RT training (Def. 2), three existing works have been proposed, corresponding to essentially two distinct approaches depending on how the perturbation risk function $\mathcal{R}_{\gamma}$ is realized. A simple augmentation-based strategy is introduced by Wang et al. (2021), referred to as corruption training, where each input $\bm{x}_{i}$ is perturbed by a sample drawn from a distribution (uniform by default) within an $\ell_{\infty}$ ball of radius $\gamma$. Robey et al. (2022) proposed a CVaR-based risk objective that emphasizes the tail of the loss distribution, thereby prioritizing samples with larger loss values. A similar work by Zhang et al. (2024a) replaces CVaR with EVaR, which leverages the entire distribution while following the same training strategy via sampling [^2] Given the conceptual similarity, categorize both EVaR and CVaR under the CVaR category in Table 1.. The “hybrid” method AT-PR targets PR while following the spirit of AT, as summarized in Table 1. It introduces a new min–max optimization objective (Eq. 20). Through multi-start PGD attacks and decision boundary exploration, AT-PR selects a local optimal AE in the norm-ball ($\bm{\delta}\in\{\bm{\delta}^{PGD}\}$) whose neighborhood constitutes the largest all-AE region $k$, i.e., $\textit{PR}(x+\delta,k)=0$.
@@ 3.3 Evaluation Metrics
In addition to clean accuracy, $\mathtt{PRBench}$ evaluates the models across three core aspects: AR, PR, and GE, cf. Table 2. AR performance is assessed by classification accuracy under adversarial attacks (Eq. 2). PR is quantified using two metrics from the literature: $\textit{PR}_{\mathcal{D}}(\gamma)$ Webb et al. (2019), the average probability that correctly classified inputs $\bm{x}_{i}$ remain correct under perturbations of radius $\gamma$ (cf. Def. 1), and $\textit{ProbAcc}(\rho)$ Robey et al. (2022), the fraction of inputs $\bm{x}$ whose $\textit{PR}(\bm{x},\gamma)$ exceeds a given threshold $1-\rho$. For the two PR metrics, $|\mathcal{D}|$ denotes the number of test samples that are correctly classified by the model, whereas for AR, $|\mathcal{D}|$ refers to the total number of test samples. GE captures robustness gap between the training and test sets. More details are deferred to Appendix D.

### v2

@@ 1 Introduction
Deep learning (DL) has demonstrated remarkable potential to drive transformative advancements across a wide range of industries, such as autonomous driving and medical diagnostics. In such safety-critical domains, robustness is a fundamental prerequisite for DL models’ pervasive deployment. Numerous studies have investigated DL robustness, leading to a range of benchmarks that systematically track the progress in the topic ling2019deepsec; guo2023comprehensive; croce2020robustbench; tang2021robustart; dong2020benchmarking; liu2025comprehensive. These efforts have predominantly focused on adversarial robustness (AR), which emphasizes the extreme worst-case scenarios by evaluating local robustness based on the existence of deterministic adversarial examples (AEs): subtle perturbations to inputs that alter model predictions szegedy2013intriguing; goodfellow2014explaining; papernot2016limitations.
@@
A more recent and practical perspective, probabilistic robustness (PR) webb2018statistical; weng2019proven; couellan2021probabilistic; baluta2021scalable; zhao_assessing_2021; tit2021efficient; robey2022probabilistically; pautov2022cc; zhang2022proa; Huang_2023_ICCV; dong2023reliability; pmlr_v206_tit23a; zhang2024prass; zhang2024protip; zhang2025adversarial; zhang2025are; zhang2023towards, employs statistical approaches to answer the question: “What is the probability that predictions remain correct under stochastic perturbations”. This probabilistic view is arguably more relevant to real-world applications than AR, as it provides an overall assessment of a model’s local robustness, accounting for scenarios where AEs may exist webb2018statistical; Huang_2023_ICCV; zhao2025probabilistic and acknowledging residual risks zhang2022proa; zhang2024prass; zhang2025adversarial that are more realistic to manage in practice.
@@
Despite its potential, most existing work on PR is restricted to evaluation, with only a few studies wang2021statistically; robey2022probabilistically; zhang2024prass; zhang2025adversarial exploring training methods specifically designed to improve PR zhao2025probabilistic. Among them, we identify three limitations. First, they use different sets of evaluation metrics, with none adopting a comprehensive set to assess the training methods holistically, making it difficult to understand which methods are truly effective/efficient and under what conditions. Second, an interesting observation is that adversarial training (AT) that primarily designed to improve AR often enhances PR as a “free by-product”. However, existing studies assess only a limited selection of AT methods, limiting the understanding of AT’s potential as a more efficient PR improvement tool. Third, there is no unified theoretical framework to compare the generalisability of PR-targeted training, making it unclear of their broader applicability in various context. These gaps and the rapid growth of PR research highlight the need for a systematic benchmark for PR training methods, which is currently missing from the literature.
@@
To bridge the gaps, we introduce $\mathtt{PRBench}$, the first benchmark dedicated to evaluating training methods for improving PR. Key features of $\mathtt{PRBench}$ include: i) A comprehensive set of metrics, covering clean accuracy, PR and AR performance, training efficiency, and generalisation error (GE); ii) A large set of training methods including most common AT methods and all PR-targeted training methods; iii) Theoretical bounds of GE are derived under a unified framework of Uniform Stability Analysis xiao2022stability; cheng2024stability to conclude comparative insights. Specifically, $\mathtt{PRBench}$ includes 222 trained models based on 7 widely adopted datasets and 10 model architectures. It uses 4 common adversarial attacks to measure AR performance, 2 types of PR metrics (with various hyperparameters, e.g., distributions of perturbations) for PR, GE metrics for both AR and PR, the clean accuracy and the training time. While the total number of training methods included is 13, $\mathtt{PRBench}$ code-base is designed to be extendable for future inclusion and comparison of new methods.
@@ 2.1 Generalisation Errors and Adversarial Robustness
AT goodfellow2014explaining; tramer2018ensemble; gowal2019scalable; madry2018towards; balunovic2020adversarial is the most common and effective empirical approach for enhancing AR. It is typically formulated as a min-max optimisation problem, where the inner max aims at finding the worst-case AE in Eq. 2, e.g., by FGSM goodfellow2014explaining and PGD madry2018towards. Then the model is trained to minimise the loss over these AEs derived from a training dataset $\mathcal{D}$:
@@ 2.2 Probabilistic Robustness
PR adopts a probabilistic view, evaluating the overall local robustness in the presence of AEs, cf. Fig. 1 (b). This probabilistic notion of robustness is acknowledged to be more practical than the worst-case AR, as many applications only require the risk of AEs to stay below an acceptable threshold, rather than eliminating them entirely webb2018statistical; zhang2022proa; zhang2024protip; zhao2025probabilistic.
@@
where $I_{\mathcal{S}}(\bm{x})$ is an indicator function that equals 1 when $\mathcal{S}$ is true and 0 otherwise; $Pr(\cdot)$ is the local distribution of inputs representing how perturbations $\bm{\delta}$ are generated, which is precisely the “input model” used by webb2018statistical; weng2019proven; zhang2024protip.
Def. 1 suggests that PR is the probability that the model prediction remains unchanged from a random perturbation $\bm{x}^{\prime}$. A “frequentist” interpretation of this expected probability is—it is the limiting relative frequency of perturbations where the output label is preserved, in an infinite sequence of independently generated perturbations zhang2024protip. In other words, the “proportion” of non-AEs in the infinite set of perturbed inputs. To evaluate PR, one of the earliest works webb2018statistical introduced a useful black-box statistical estimator, especially for cases where PR is very high. Later, more efficient white-box statistical estimators were proposed in pmlr_v206_tit23a. Additionally, PR has been extended to applications such as explainable AI Huang_2023_ICCV and text-to-image models zhang2024protip. For an overview of PR estimators, we refer readers to zhao2025probabilistic.
In contrast to the relatively extensive studies on PR estimators, research on training methods for PR remains scarce. To the best of our knowledge, four studies wang2021statistically; robey2022probabilistically; zhang2024prass; zhang2025adversarial explicitly motivated to develop training methods for improving PR. The first three can be categorized as Risk-based Training (RT) methods, which shift away from the worst-case perspective towards a training paradigm based on statistical risks induced by distributional perturbations, incorporating perturbation risk functions into the objective. We formally introduce its general formulation as:
@@
The specific choice of $\mathcal{R}_{\gamma}$ in RT methods varies: wang2021statistically uses the identity function, while robey2022probabilistically; zhang2024prass adopt functions calculating the Conditional Value-at-Risk (CVaR) and Entropic Value-at-Risk (EVaR), respectively. Essentially, rather than training on optimized AEs like AT methods for AR, RT methods train on statistical risks quantified by sampling stochastic perturbations from the perturbation distribution.
A very recent idea of adapting AT for PR was proposed in zhang2025adversarial, which does not follow RT paradigm of Def. 2. Instead, their approach aligns with the traditional AT formulation in Eq. 3. Through multi-start PGD attacks and decision boundary exploration, AT-PR identifies an optimal AE whose neighborhood constitutes the largest all-AE region. Intuitively, training over such an optimal AE would reduce the overall “proportion” of AEs in the local norm ball and thus improve PR. We classify it as “hybrid” given it is targeting PR but adapting AT by solving a new min-max problem. Both RT methods and the “hybrid” method AT-PR constitute the current class of PR-targeted method. Further details for all four training methods are provided in Appendix C.
@@ 2.3 Robust Overfitting
While AT is regarded as one of the most promising methods for enhancing AR, empirical studies rice2020overfitting; gowal2020uncovering have shown that it suffers significantly from overfitting. To understand this phenomenon, several theoretical studies have been conducted under different frameworks, including VC-dimension montasser2019vc; attias2022improved, Rademacher complexity khim2018adversarial; yin2019rademacher; awasthi2020adversarial; xiao2022adversarial, and Uniform Algorithmic Stability farnia2021train; xing2021algorithmic; xiao2022adaptive. Among them, one of the most notable contributions is by xiao2022stability, which introduces the concept of $\eta$-approximate $\beta$-smoothness, as defined in Eq. 6. This work xiao2022stability analyses the surrogate loss of the max function and show that generalisation is affected by an additional term proportional to $\eta$.
Definition 3 (Approximate Smoothness xiao2022stability)
@@
This assumption has been widely adopted and validated in prior work farnia2021train; xing2021algorithmic; xiao2022adaptive; xiao2022stability, where it serves as a foundational premise for analyzing algorithmic stability and generalisation under AT. We show that the GE is bounded in Thm. 1.
@@
Thm. 1 follows from the results in xiao2022stability, where the authors conclude that the additional term $n\varphi\phi$ contributes to robust overfitting. The complete proof corresponds to the special case $\lambda=0$ of Thm. 3 in Appendix E.2. And the proof for uniform stability is shown in Appendix E.5.
@@ 3 Description of PRBench
All existing robustness benchmarks focus exclusively on AR guo2023comprehensive; croce2020robustbench; tang2021robustart; dong2020benchmarking; liu2025comprehensive, with more related benchmarks summarized in Appendix A.5. In contrast, our $\mathtt{PRBench}$ is the first benchmark dedicated to PR. This section describes the models, datasets, training methods, and evaluation metrics in $\mathtt{PRBench}$.
@@ 3.2 Training Methods
For AT, we consider the following representative methods. The PGD attack is widely regarded as the standard approach, which minimizes the adversarial cross-entropy loss $\mathcal{L}_{\textit{CE}}$. Building on this, logit pairing methods kannan2018adversarial, including adversarial logit pairing (ALP) and clean logit pairing (CLP), augment the objective with a regularization term coupling natural and AEs. TRADES further introduces a KL-based regularization, defining its objective as a combination of the natural loss and the KL divergence $\mathcal{L}_{\textit{KL}}$ between predictions on clean and adversarial inputs. MART follows the same KL-based regularization but augments it with the adversarial cross-entropy loss $\mathcal{L}_{\textit{CE}}$, while additionally emphasizing misclassified examples through larger penalty weights. Overall, existing AT methods can be characterized by different uses of $\mathcal{L}_{\textit{CE}}$ and $\mathcal{L}_{\textit{KL}}$, either individually or in combination, for both AE generation and training. We identify a missing configuration and introduce KL-PGD, a combination that generates AEs using $\mathcal{L}_{\textit{KL}}$ (as in TRADES) and trains the model with $\mathcal{L}_{\textit{CE}}$ (as in PGD). KL-PGD is designed as a diagnostic variant to disentangle the effects of loss functions from perturbation generation strategies, thereby providing a controlled comparison point for systematically evaluating the relative contributions of different optimization objectives.
@@
By instantiating the general formulation of RT training (Def. 2), three existing works have been proposed, corresponding to essentially two distinct approaches depending on how the perturbation risk function $\mathcal{R}_{\gamma}$ is realized. A simple augmentation-based strategy is introduced by wang2021statistically, referred to as corruption training, where each input $\bm{x}_{i}$ is perturbed by a sample drawn from a distribution (uniform by default) within an $\ell_{\infty}$ ball of radius $\gamma$. robey2022probabilistically proposed a CVaR-based risk objective that emphasizes the tail of the loss distribution, thereby prioritizing samples with larger loss values. A similar work by zhang2024prass replaces CVaR with EVaR, which leverages the entire distribution while following the same training strategy via sampling [^2] Given the conceptual similarity, categorize both EVaR and CVaR under the CVaR category in Table 1.. The “hybrid” method AT-PR targets PR while following the spirit of AT, as summarized in Table 1. It introduces a new min–max optimization objective (Eq. 20). Through multi-start PGD attacks and decision boundary exploration, AT-PR selects a local optimal AE in the norm-ball ($\bm{\delta}\in\{\bm{\delta}^{PGD}\}$) whose neighborhood constitutes the largest all-AE region $k$, i.e., $\textit{PR}(x+\delta,k)=0$.
@@ 3.3 Evaluation Metrics
In addition to clean accuracy, $\mathtt{PRBench}$ evaluates the models across three core aspects: AR, PR, and GE, cf. Table 2. AR performance is assessed by classification accuracy under adversarial attacks (Eq. 2). PR is quantified using two metrics from the literature: $\textit{PR}_{\mathcal{D}}(\gamma)$ webb2018statistical, the average probability that correctly classified inputs $\bm{x}_{i}$ remain correct under perturbations of radius $\gamma$ (cf. Def. 1), and $\textit{ProbAcc}(\rho)$ robey2022probabilistically, the fraction of inputs $\bm{x}$ whose $\textit{PR}(\bm{x},\gamma)$ exceeds a given threshold $1-\rho$. For the two PR metrics, $|\mathcal{D}|$ denotes the number of test samples that are correctly classified by the model, whereas for AR, $|\mathcal{D}|$ refers to the total number of test samples. GE captures robustness gap between the training and test sets. More details are deferred to Appendix D.